OpenStudy (anonymous):

Last question... f(x) = x + 2 and g(x) = x2 – 6x + 3. Find g(f(x))

OpenStudy (mertsj):

g(f(x)) means replace each x in the function g with the function f so: g(f(x)) = (x+2)^2 - 6(x+2) +3 g(f(x)) = x^2 +4x +4 -6x -12 + 3 g(f(x)) = x^2 -2x -5

OpenStudy (mertsj):

Now I am going to double check my work.

OpenStudy (anonymous):

lol..I think your right

OpenStudy (mertsj):

Yep, I think so.

OpenStudy (wasiqss):

write mr mertsj, im hea to monitor u

OpenStudy (mertsj):

mr

OpenStudy (wasiqss):

:D